# On the PM2.5 – Mortality Relationship: A Bayesian Dynamic Model for Spatio-Temporal Confounding

### v1

@@ Acknowledgements
Hersbach et al. (2018) was downloaded from the C3S (2023). INERIS et al. (2022) was downloaded from the CAMS (2023). The results contain modified C3S and CAMS information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The research of Zaccardi, Valentini and Ippoliti is funded by the European Union - NextGenerationEU, research project PRIN2022 CoEnv - Complex Environmental Data and Modeling (CUP 2022E3RY23). Schmidt is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999).
@@ Appendix A Data Collection
  • Exposure data. Data on PM2.5 concentrations (expressed in $\mu g/m^{3}$) are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). Hourly data are available at a spatial resolution of $0.1^{\circ}\times 0.1^{\circ}$, so we use zonal statistics to aggregate the estimates to the HD level, and then average across time to obtain daily data. To evaluate the short-term effect of interest, a lag of one day is applied to the exposure’s time series.
@@ Appendix B Descriptive Statistics
Appendix C Details on Competing Models
@@
$\vartheta_{it}$ $=o_{it}+\mathbf{A}_{it}^{\prime}\mathbf{a}_{i}+\varepsilon_{it},\quad\varepsilon_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{\varepsilon}^{2})$
@@
where $\mathbf{A}_{it}$ is a $p_{a}$-dimensional vector of covariates (i.e., intercept, exposure and measured confounders) and $\mathbf{a}_{i}$ is the corresponding vector of coefficients. The vector $\bm{\mu}_{a}$ is of main interest, specifically its element(s) corresponding to the overall exposure effect.
The specification of $\mathbf{A}_{it}$ depends on the approach under consideration as listed below.
  • Null: the model that does not account for unmeasured confounding, for which $\mathbf{A}_{it}=(1,X_{it},{{}^{m}{\bf Z}}_{it}^{\prime})^{\prime}$. The regression coefficients are allowed to vary across areas but not in time.
  • GLMadj: the model that includes smooth function of time, $h(t;df)$, modelled as a natural cubic spline with $df$ degrees of freedom. Let $\mathbf{B}_{t}$ denote the vector of basis functions evaluated at time $t$. Then, we define $\mathbf{A}_{it}=(1,X_{it},{{}^{m}{\bf Z}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
@@
    Then, $\mathbf{A}_{it}=(1,X_{it}D_{1t},\dots,X_{it}D_{4t},{{}^{m}{\bf Z}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$. Henceforth, the Dummy model is similar to GLMadj, but it also captures season-specific exposure effects.
@@
  • Periodic: the exposure effect is assumed to be a linear combination of trigonometric functions, as in Peng et al. (2005), Chen et al. (2013). In this case,
    $\mathbf{A}_{it}=(1,X_{it},X_{it}\cos(2\pi t/365),X_{it}\sin(2\pi t/365),{{}^{m}{\bf Z}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
@@
  • JDZ: the model proposed by Janes et al. (2007), where the local exposure effect is of interest, hence $\mathbf{A}_{it}=(1,X_{it}-\widehat{X_{t}},\widehat{X_{t}},{{}^{m}{\bf Z}}_{it}^{\prime},\mathbf{B}_{t}^{*\prime})^{\prime}$. Here, $\widehat{X_{t}}$ denotes the overall temporal trend (see also Section 3), and $\mathbf{B}_{t}^{*}$ denotes the vector of basis functions from a natural cubic spline with $df-1$ degrees of freedom, evaluated at time $t$.
@@
The unknown function $h(\cdot;\cdot)$ is modelled as a natural cubic spline of time with 8 degrees of freedom per year (so $df=40$), and is used to control for unmeasured confounding. The overall trend, $\widehat{X_{t}}$, is obtained as the fitted values of a regression of $X_{it}$ on a natural cubic spline of time with $4$ degrees of freedom per year. Bases from natural cubic splines of temperature (with 7 degrees of freedom) and of relative humidity (with 4 degrees of freedom) are included as measured confounders. ${{}^{m}{\bf Z}}_{it}$ also includes the indicator variables for the day of the week, for holidays, and for the lockdown periods.
Appendix D Posterior Full Conditional Distributions for the Proposed DGLM
$\bm{\vartheta}=\mathbf{o}+\mathbf{G}\bm{\gamma}+\tilde{\mathbf{u}}\,,$
where $\bm{\vartheta}=(\bm{\vartheta}_{1}^{\prime},\dots,\bm{\vartheta}_{T}^{\prime})^{\prime}$, $\mathbf{o}=(\mathbf{o}_{1}^{\prime},\dots,\mathbf{o}_{T}^{\prime})^{\prime}$, $\tilde{\mathbf{u}}=(\tilde{\mathbf{u}}_{1}^{\prime},\dots,\tilde{\mathbf{u}}_{T}^{\prime})^{\prime}$, $\bm{\gamma}=(\tilde{\bm{\beta}}_{1}^{\prime},\tilde{\bm{\beta}}_{2}^{\prime},\dots,\tilde{\bm{\beta}}_{{T}}^{\prime},\bm{\alpha}_{m}^{\prime})^{\prime}$, and
$\mathbf{G}=\begin{bmatrix}\mathbf{\tilde{X}}_{\,\bm{\cdot}1}&&&{}^{m}\mathbf{Z}_{1}\\ &\ddots&&\vdots\\ &&\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}}&{}^{m}\mathbf{Z}_{T}\end{bmatrix}\,.$
Posterior samples are obtained using a Gibbs sampler with the following full conditionals:
$\bm{\gamma}|\bullet\sim\mathcal{N}(\overline{\mathbf{V}}\overline{\bm{\mu}},\overline{\mathbf{V}})$
$\sigma_{w1}^{2}|\bullet\sim\mathcal{IG}\left(a_{w1}+\frac{{T}-1}{2},b_{w1}+\frac{1}{2}\sum_{t=2}^{{T}}(\beta_{1t}-\beta_{1,t-1})^{2}\right)$
$\bm{\varSigma}_{w0}|\bullet\sim\mathcal{IW}\left(\mathbf{a}_{w0}+{T}-1,\mathbf{S}_{w0}+\sum_{t=2}^{{T}}(\tilde{\bm{\beta}}_{0t}-\tilde{\bm{\beta}}_{0,t-1})(\tilde{\bm{\beta}}_{0t}-\tilde{\bm{\beta}}_{0,t-1})^{\prime}\right)$
$\vartheta_{it}|\bullet\propto\exp\left\{Y_{it}\vartheta_{it}-\exp(\vartheta_{it})-\frac{1}{2\tau_{u}^{2}}(\vartheta_{it}-o_{it}-\vartheta_{it}^{*})^{2}\right\}\quad\mbox{for}\quad i=1,\dots,{N}\quad\mbox{and}\quad t=1,\dots,{T}$
$\tilde{\tau}_{u}^{2}|\bullet\sim\mathcal{IG}\left(a_{u}+\frac{{N}{T}}{2},b_{u}+\frac{1}{2}\sum_{t=1}^{{T}}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})^{\prime}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})\right)$
where $\vartheta_{it}^{*}$ is a generic element of $\bm{\vartheta}_{t}^{*}=(\vartheta_{11}^{*},\vartheta_{21}^{*},\dots,\vartheta_{{N}{T}}^{*})^{\prime}=\mathbf{G}\bm{\gamma}$, and the symbol $\bullet$ denotes data and all the parameters except for the parameter that is being updated. For $\vartheta_{it}$, $i=1,\dots,{N}$, $t=1,\dots,{T}$, we use Metropolis-Hastings steps with a Gaussian proposal distribution.
For the posterior full conditional of $\bm{\gamma}$, two more matrices need to be defined:
$\mathbf{H}=\begin{bmatrix}\mathbf{I}_{N}&&&&&\\ -\mathbf{I}_{N}&\mathbf{I}_{N}&&&&\\ &-\mathbf{I}_{N}&\mathbf{I}_{N}&&&\\ &&\ddots&\ddots&&\\ &&&-\mathbf{I}_{N}&\mathbf{I}_{N}&\\ &&&&&\mathbf{I}_{p_{m}}\end{bmatrix}\quad\mbox{and}\quad\mathbf{S}=\begin{bmatrix}\mathbf{V}_{\beta}&&&&&\\ &\bm{\varSigma}_{w}&&&&\\ &&\ddots&&&\\ &&&\bm{\varSigma}_{w}&&\\ &&&&\bm{\varSigma}_{w}&\\ &&&&&10^{6}\mathbf{I}_{{p_{m}}}\end{bmatrix}\,.$
Hence, this posterior full conditional is multivariate normal, $\bm{\gamma}|\bullet\sim\mathcal{N}(\overline{\mathbf{V}}\overline{\bm{\mu}},\overline{\mathbf{V}})$, with
$\overline{\mathbf{V}}=(\mathbf{K}+\mathbf{G}^{\prime}\mathbf{G}/\tilde{\tau}_{u}^{2})^{-1}\quad\mbox{and}\quad\overline{\bm{\mu}}=\mathbf{G}^{\prime}\bm{\vartheta}/\tilde{\tau}_{u}^{2}\,,$
where $\mathbf{K}=\mathbf{H}^{\prime}\mathbf{S}^{-1}\mathbf{H}$. We refer to Chan and Jeliazkov (2009) for further computational details.
@@ E.1 Data Generation Mechanism
To simulate the variables, we use the following data generation mechanism. The spatial domain comprises ${N}=100$ contiguous HDs, while the temporal dimension is set to ${T}=300$. Let ${\bf X}=({\bf X}_{1},\dots,{\bf X}_{{T}})^{\prime}$, with ${\bf X}_{t}=(X_{1t},\dots,X_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the exposure, and let ${\bf Z}=({\bf Z}_{1},\dots,{\bf Z}_{{T}})^{\prime}$, with ${\bf Z}_{t}=(Z_{1t},\dots,Z_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the unmeasured confounder. It is thus assumed that
@@
where $\bm{\mu}_{x}$ and $\bm{\mu}_{z}$ are the spatio-temporal means generated from Fourier basis expansions such that their correlation is approximately $0.5$. It is assumed that the covariance matrices can be factorized as $\bm{\varSigma}_{x}=\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{x}^{1/2\prime}$, $\bm{\varSigma}_{z}=\bm{\varSigma}_{z}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, $\bm{\varSigma}_{xz}=\bm{\varSigma}_{zx}^{\prime}=\rho_{xz}\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, and $\rho_{xz}={\rm\mathbb{C}or}(X_{it},Z_{it})$, for $i=1,\dots,{N}$, and $t=1,\dots,{T}$.
@@
$\bm{\Omega}_{x,pcar}={\bf D}_{w}-\phi_{x}^{S}{\bf W}\,,$ (10)
where ${\bf D}_{w}$ is a diagonal matrix with the number of neighbors of each HD on the diagonal, ${\bf W}$ is the binary adjacency matrix, and $\phi_{x}^{S}\in(0,1)$ is the spatial autocorrelation parameter. ${\bf W}$ is an ${N}\times{N}$ matrix with entries $w_{ij}=1$ if areal units $i$ and $j$ share a border, and $w_{ij}=0$ otherwise; also, $w_{ii}=0$.
@@
$\vartheta_{it}$ $=o_{it}+\beta_{0i}+\beta_{1t}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})\,,$ (12)
where $\mu_{it}=\exp(\vartheta_{it})$, $o_{it}=\log(E_{it})$ is the offset term, and $\tau_{u}^{2}=10^{-5}$. The expected counts, $E_{it}$, are obtained by using the resident population aged $\geq 65$ years and the age-specific standardized mortality rates of year 2018. The population and the rates are assumed to be constant in time.

### v2

@@ Acknowledgements
Hersbach et al. (2018) was downloaded from the C3S (2023). INERIS et al. (2022) was downloaded from the CAMS (2023). The results contain modified C3S and CAMS information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains. The research of C.Z., P.V., and L.I. is funded by the European Union - NextGenerationEU. C.Z.: PRIN PNRR 2022 (Bando 2023) SLIDE - Stochastic Modeling of Compound Events (P2022KZJTZ - CUP D53D23018920001). P.V. and L.I.: research project PRIN2022 CoEnv - Complex Environmental Data and Modeling (2022E3RY23 - CUP D53D23011080006). A.M.S. is grateful for financial support from the Natural Sciences and Engineering Research Council (NSERC) of Canada (Discovery Grant RGPIN-2017-04999).
@@ Appendix A Data Collection
  • Exposure data. Data on PM2.5 concentrations (expressed in $\mu g/m^{3}$) are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). Hourly data are available at a spatial resolution of $0.1^{\circ}\times 0.1^{\circ}$, so we use zonal statistics to aggregate the estimates to the HD level, and then average across time to obtain daily data.
@@ Appendix B Descriptive Statistics
Appendix C Posterior Full Conditional Distributions for the Proposed DGLM
$\bm{\vartheta}=\mathbf{o}+\mathbf{G}_{1}\bm{\gamma}_{1}+\mathbf{G}_{2}\bm{\gamma}_{2}+\mathbf{G}_{3}\bm{\gamma}_{3}+\mathbf{G}_{4}\bm{\gamma}_{4}+\tilde{\mathbf{u}}\,,$
where $\bm{\vartheta}=(\bm{\vartheta}_{1}^{\prime},\dots,\bm{\vartheta}_{T}^{\prime})^{\prime}$, $\mathbf{o}=(\mathbf{o}_{1}^{\prime},\dots,\mathbf{o}_{T}^{\prime})^{\prime}$, $\tilde{\mathbf{u}}=(\tilde{\mathbf{u}}_{1}^{\prime},\dots,\tilde{\mathbf{u}}_{T}^{\prime})^{\prime}$, $\bm{\gamma}_{1}=(\bm{\delta}_{01}^{*\prime},\bm{\delta}_{02}^{*\prime},\dots,\bm{\delta}_{0{T}}^{*\prime})^{\prime}$, $\bm{\gamma}_{2}=(\delta_{01},\delta_{11},\dots,\delta_{0{T}},\delta_{1{T}})^{\prime}$, $\bm{\gamma}_{3}=(\tilde{\bm{\delta}}_{0}^{\prime},\tilde{\bm{\delta}}_{1}^{\prime})^{\prime}$, $\bm{\gamma}_{4}=(\overline{\delta}_{0},\overline{\delta}_{1},\bm{\alpha}^{\prime})^{\prime}$. The observation matrices are $\mathbf{G}_{1}={\bf I}_{{N}{T}}$,
$\mathbf{G}_{2}=\begin{bmatrix}{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}1}&&&\\ &&{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}2}&&\\ &&&\ddots&\ddots&\\ &&&&{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}}\end{bmatrix}\,,\qquad\mathbf{G}_{3}=\begin{bmatrix}{\bf I}_{N}&\mbox{diag}(\mathbf{\tilde{X}}_{\,\bm{\cdot}1})\\ \vdots&\vdots\\ {\bf I}_{N}&\mbox{diag}(\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}})\end{bmatrix}\,,\qquad\mathbf{G}_{4}=\begin{bmatrix}{\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}1}&{\mathbf{M}}_{1}\\ \vdots&\vdots&\vdots\\ {\bf 1}_{N}&\mathbf{\tilde{X}}_{\,\bm{\cdot}{T}}&{\mathbf{M}}_{T}\end{bmatrix}\,.$
Posterior samples are obtained using a Gibbs sampler with the following full conditionals, where the symbol $\bullet$” denotes data and all the parameters except for the parameter that is being updated.
  1. 1.
    Sample regression coefficients, $\bm{\gamma}_{j}$, for $j=1,2,3,4$:
    $\bm{\gamma}_{j}|\bullet\sim\mathcal{N}(\overline{\mathbf{V}}_{j}\overline{\bm{\mu}}_{j},\overline{\mathbf{V}}_{j})\,,$
    where $\overline{\mathbf{V}}_{j}=(\mathbf{K}_{j}+\mathbf{G}_{j}^{\prime}\mathbf{G}_{j}/\tilde{\tau}_{u}^{2})^{-1}$ and $\overline{\bm{\mu}}_{j}=\mathbf{G}_{j}^{\prime}\bm{\eta}_{j}/\tilde{\tau}_{u}^{2}$, where $\bm{\eta}_{j}=\bm{\vartheta}-\mathbf{o}-\sum_{l\neq j}\mathbf{G}_{l}\bm{\gamma}_{l}$, $\mathbf{K}_{j}=\mathbf{H}_{j}^{\prime}\mathbf{S}_{j}^{-1}\mathbf{H}_{j}$, and
    $\mathbf{H}_{1}=\begin{bmatrix}\mathbf{I}_{N}&&&\\ -\mathbf{I}_{N}&\mathbf{I}_{N}&\\ &\ddots&\ddots&\\ &&-\mathbf{I}_{N}&\mathbf{I}_{N}\\ \end{bmatrix}\quad\mbox{and}\quad\mathbf{S}_{1}^{-1}=\begin{bmatrix}\mathbf{V}_{\delta}^{*-1}&&&\\ &\sigma_{w0}^{*-2}(\mathbf{D-W})&&\\ &&\ddots&\\ &&&\sigma_{w0}^{*-2}(\mathbf{D-W})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{2}=\begin{bmatrix}\mathbf{I}_{2}&&&\\ -\mathbf{I}_{2}&\mathbf{I}_{2}&\\ &\ddots&\ddots&\\ &&-\mathbf{I}_{2}&\mathbf{I}_{2}\end{bmatrix}\quad\mbox{and}\quad\mathbf{S}_{2}^{-1}=\begin{bmatrix}\mbox{diag}(V_{\delta 0}^{-1},V_{\delta 1}^{-1})&&&\\ &\mbox{diag}(\sigma_{w0}^{-2},\sigma_{w1}^{-2})&&\\ &&\ddots&\\ &&&\mbox{diag}(\sigma_{w0}^{-2},\sigma_{w1}^{-2})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{3}={\bf I}_{2{N}}\quad\mbox{and}\quad\mathbf{S}_{3}^{-1}=\begin{bmatrix}\sigma_{\delta 0}^{-2}(\mathbf{D-W})&\\ &\sigma_{\delta 1}^{-2}(\mathbf{D-W})\\ \end{bmatrix}\,,$
    $\mathbf{H}_{4}={\bf I}_{{p_{m}}+2}\quad\mbox{and}\quad\mathbf{S}_{4}^{-1}=10^{-6}{\bf I}_{{p_{m}}+2}\,.$
  2. 2.
    Sample variance parameters of temporal effects, for $k=0,1$:
    $\sigma_{wk}^{2}|\bullet\sim\mathcal{IG}\left(a_{wk}+\frac{{T}-1}{2},b_{wk}+\frac{1}{2}\sum_{t=2}^{{T}}(\delta_{kt}-\delta_{k,t-1})^{2}\right)$
  3. 3.
    Sample variance parameters of spatial effects, for $k=0,1$:
    $\sigma_{\delta k}^{2}|\bullet\sim\mathcal{IG}\left(a_{\delta k}+\frac{{N}-1}{2},b_{\delta k}+\frac{1}{2}\tilde{\bm{\delta}}_{k}^{\prime}(\mathbf{D}-\mathbf{W})\tilde{\bm{\delta}}_{k}\right)$
  4. 4.
    Sample variance parameter of spatio-temporal effect:
    $\sigma_{w0}^{*2}|\bullet\sim\mathcal{IG}\left(a_{w0}^{*}+\frac{({N}-1)({T}-1)}{2},b_{w0}^{*}+\frac{1}{2}\sum_{t=2}^{{T}}(\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*})^{\prime}(\mathbf{D}-\mathbf{W})(\bm{\delta}_{0t}^{*}-\bm{\delta}_{0,t-1}^{*})\right)$
  5. 5.
    Sample the latent variables, $\vartheta_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$, using Metropolis-Hastings steps with a Gaussian proposal distribution:
    $\vartheta_{it}|\bullet\propto\exp\left\{Y_{it}\vartheta_{it}-\exp(\vartheta_{it})-\frac{1}{2\tilde{\tau}_{u}^{2}}(\vartheta_{it}-o_{it}-\vartheta_{it}^{*})^{2}\right\}\,,$
    where $\vartheta_{it}^{*}$ is a generic element of $\bm{\vartheta}_{t}^{*}=(\vartheta_{11}^{*},\vartheta_{21}^{*},\dots,\vartheta_{{N}{T}}^{*})^{\prime}={\bf 1}_{N}(\overline{\delta}_{0}+\delta_{0t})+\tilde{\bm{\delta}}_{0}+\bm{\delta}_{0t}^{*}+\mathbf{\tilde{X}}_{t}(\overline{\delta}_{1}+\delta_{1t})+\mathbf{\tilde{X}}_{t}\odot\tilde{\bm{\delta}}_{1}+{\mathbf{M}}_{t}\bm{\alpha}$.
  6. 6.
    Sample the measurement error variance:
    $\tilde{\tau}_{u}^{2}|\bullet\sim\mathcal{IG}\left(a_{u}+\frac{{N}{T}}{2},b_{u}+\frac{1}{2}\sum_{t=1}^{{T}}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})^{\prime}(\bm{\vartheta}_{t}-\mathbf{o}_{t}-\bm{\vartheta}_{t}^{*})\right)\,.$
Appendix D Details on Competing Models
@@
$\vartheta_{it}$ $=o_{it}+\mathbf{A}_{it}^{\prime}\mathbf{a}_{i}+\mathbf{C}_{it}^{\prime}\mathbf{c}_{i}+\varepsilon_{it},\quad\varepsilon_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{\varepsilon}^{2})$
@@
where $\mathbf{A}_{it}$ is a $p_{a}$-dimensional vector of covariates related to the exposure effect, $\mathbf{C}_{it}$ is a $p_{c}$-dimensional vector of covariates (i.e., intercept and measured confounders), and $\mathbf{a}_{i}$ and $\mathbf{c}_{i}$ are the corresponding vectors of coefficients. The vector $\bm{\mu}_{a}$ and the matrix $\bm{\varSigma}_{a}$ are of main interest since they correspond to the overall exposure effect and the between-site variability, respectively.
The specification of $\mathbf{A}_{it}$ and $\mathbf{C}_{it}$ depends on the approach under consideration as listed below.
  • Null: the model that does not account for unmeasured confounding, for which $\mathbf{A}_{it}=X_{it}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime})^{\prime}$. The regression coefficients are allowed to vary across areas but not in time.
  • GLMadj: the model that includes smooth function of time, $h(t;df)$, modelled as a natural cubic spline with $df$ degrees of freedom. Let $\mathbf{B}_{t}$ denote the vector of basis functions evaluated at time $t$. Then, we define $\mathbf{A}_{it}=X_{it}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
@@
    Then, $\mathbf{A}_{it}=(X_{it}D_{1t},\dots,X_{it}D_{4t})^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$. Henceforth, the Dummy model is similar to GLMadj, but it also captures season-specific exposure effects.
@@
  • Periodic: the exposure effect is assumed to be a linear combination of trigonometric functions, as in Peng et al. (2005), Chen et al. (2013). In this case, $\mathbf{A}_{it}=(X_{it},X_{it}\cos(2\pi t/365),X_{it}\sin(2\pi t/365))^{\prime}$ and $\mathbf{C}_{it}=(1,{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{\prime})^{\prime}$.
@@
  • JDZ: the model proposed by Janes et al. (2007), where the local exposure effect is of interest, hence $\mathbf{A}_{it}=(X_{it}-\widehat{X}_{it})^{\prime}$ and $\mathbf{C}_{it}=(1,\widehat{X}_{it},{{\bf M}}_{it}^{\prime},\mathbf{B}_{t}^{*\prime})^{\prime}$. Here, $\widehat{X}_{it}$ denotes the overall temporal trend (see also Section 3), and $\mathbf{B}_{t}^{*}$ denotes the vector of basis functions from a natural cubic spline with $df-1$ degrees of freedom, evaluated at time $t$.
@@
The unknown function $h(\cdot;\cdot)$ is modelled as a natural cubic spline of time with $df=43$ degrees of freedom (i.e., 8 equally-spaced knots per year plus 3 more knots to account for the peaks in the SMR time series), and is used to control for unmeasured confounding. The overall trend, $\widehat{X}_{t}$, is obtained as the fitted values of a regression of $X_{it}$ on a natural cubic spline of time with $4$ degrees of freedom per year. In the real data application, bases from natural cubic splines of temperature (with 7 degrees of freedom) and of relative humidity (with 4 degrees of freedom) are included as measured confounders. ${{\bf M}}_{it}$ also includes the indicator variables for the day of the week, for holidays, and for the lockdown periods.
@@ E.1 Data Generation Mechanism
To simulate the variables, we use the following data generation mechanism. The spatial domain comprises ${N}=117$ contiguous HDs, while the temporal dimension is set to ${T}=600$. Let ${\bf X}=({\bf X}_{1},\dots,{\bf X}_{{T}})^{\prime}$, with ${\bf X}_{t}=(X_{1t},\dots,X_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the exposure, and let ${\bf Z}=({\bf Z}_{1},\dots,{\bf Z}_{{T}})^{\prime}$, with ${\bf Z}_{t}=(Z_{1t},\dots,Z_{{N}t})^{\prime}$, $t=1,\dots,{T}$, be the unmeasured confounder. It is thus assumed that
@@
where $\bm{\mu}_{x}$ and $\bm{\mu}_{z}$ are the spatio-temporal means generated from Fourier basis expansions. It is assumed that the covariance matrices can be factorized as $\bm{\varSigma}_{x}=\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{x}^{1/2\prime}$, $\bm{\varSigma}_{z}=\bm{\varSigma}_{z}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, $\bm{\varSigma}_{xz}=\bm{\varSigma}_{zx}^{\prime}=\rho_{xz}\bm{\varSigma}_{x}^{1/2}\bm{\varSigma}_{z}^{1/2\prime}$, and $\rho_{xz}={\rm\mathbb{C}or}(\mathbf{X},\mathbf{Z})$.
@@
$\bm{\Omega}_{x,pcar}={\bf D}-\phi_{x}^{S}{\bf W}\,,$ (12)
where ${\bf D}$ is a diagonal matrix with the number of neighbors of each HD on the diagonal, ${\bf W}$ is the binary adjacency matrix, and $\phi_{x}^{S}\in(0,1)$ is the spatial autocorrelation parameter. ${\bf W}$ is an ${N}\times{N}$ matrix with entries $w_{ij}=1$ if areal units $i$ and $j$ share a border, and $w_{ij}=0$ otherwise; also, $w_{ii}=0$.
@@
$\vartheta_{it}$ $={\color[rgb]{0,0,0}o_{it}+\beta_{0}+\beta_{1it}X_{it}+Z_{it}+u_{it},\quad u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})}\,,$ (14)
where $\mu_{it}=\exp(\vartheta_{it})$, $o_{it}=\log(E_{it})$ is the offset term, and $\tau_{u}^{2}=10^{-3}$. The expected counts, $E_{it}$, are obtained by using the resident population aged $\geq 65$ years and the age-specific standardized mortality rates of year 2018. The population and the rates are assumed to be constant in time.